OpenStudy (anonymous):

I REAAAAAALLLLLLY NEED HELP! :'( express cos80ºcos70º+sin80ºsin70º as a single trigonometric function of a positive acute angle.

OpenStudy (anonymous):

well the answer is;0.86255518

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

you need to use the double angle formula for cosine: \[\cos(\theta-\phi)=\cos(\theta)\cos(\phi)-\sin(\theta)\sin(\phi)\]

OpenStudy (anonymous):

You have \[\theta = 80, \phi = 70\]

OpenStudy (anonymous):

Correction - the sign on the right-hand side should be +, not -.

OpenStudy (anonymous):

your late dude

OpenStudy (anonymous):

Not really - he asked for an explanation.

OpenStudy (anonymous):

but cool cause my friend needed the formula thanks

OpenStudy (anonymous):

or she...